OpenStudy (anonymous):

if x represents the total number of bags a person checks and each person checks more than two bags, the equation below can be used to determine C, the total fee for all the bags. 20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all?

jimthompson5910 (jim_thompson5910):

20 + 25 + 50(x - 2) = C 20 + 25 + 50(x - 2) = 145 45 + 50(x - 2) = 145 45 + 50x - 100 = 145 50x - 55 = 145 50x = 145+55 50x = 200 x = 200/50 x = 4 So she checked 4 bags.